What is the equation of the line, in standard form, that passes through (4, -3) and is parallel to the line whose equation is 4x + y - 2 = 0?
Found 3 solutions by mananth, ikleyn, MathTherapy:
Answer by mananth(16949)

4x+y=2
Find the slope of this line
y=-4x+2
y=-4x+2
Compare this equation with y=mx+b
slope m = -4
The slope of a line parallel to the above line will be the same
The slope of the required line will be -4
m=-4,point(4,-3)
Find b by plugging the values of m & the point in
y=mx+b
-3=-16.00+b
b=13
m=-4
Plug value of the slope and b
The required equation is y=-4x+13

Answer by ikleyn(53742)

.
What is the equation of the line, in standard form, that passes through (4, -3) and is parallel to the line
whose equation is 4x + y - 2 = 0?
~~~~~~~~~~~~~~~~~~~~~~~~


        It can be solved in much more efficient way than @mananth does it in his post.


The given line equation is 4x + y - 2 = 0.

It is the same as

    4x + y = 2.    (1)


Any parallel line has the form

    4x + y = c      (2)

with the same form '4x + y' in the left side and some constant 'c' in the right side.


To find 'c', we substitute coordinates of the given point (4,-3) into equation (2).
We get 

    4*4 + (-3) = 16 - 3 = 13 = c.


Thus the sough equation is

    4x + y = 13.


It is the "standard form" line equation.



ANSWER.  The "standard form" line equation is  4x + y = 13.

Solved.

----------------------------

Two post-solution notices


        (1)   Doing this way,  you should not worry about the slope.

        (2)   The final equation in the @mananth post is  NOT  a standard form.



Answer by MathTherapy(10801)

What is the equation of the line, in standard form, that passes through (4, -3) and is parallel to the line whose equation is 4x + y - 2 = 0?

Equation that sought-equation is parallel to: 4x + y - 2 = 0, or in STANDARD form: 4x + y = 2
As the lines are parallel, left-side of sought-equation is the same.
We then get: 4x + y = 4(x) + y
             4x + y = 4(4) + - 3 ----- Substituting given point, (4, - 3), on right-side
             4x + y = 16 - 3
           4x + y = 13 <=== Sought-equation, in STANDARD FORM